OpenStudy (anonymous):

I need help getting the solution set of |x-6|=4

OpenStudy (anonymous):

|x-6|=4 is same as (x-6)=4 or x-6=-4

OpenStudy (anonymous):

are you trying to find x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I came up with x=10 or x=2 but I'm not sure that is correct.

OpenStudy (anonymous):

x-6=4 x=4+6 x=10

OpenStudy (anonymous):

you would be correct

OpenStudy (anonymous):

|x-2|+4=3

OpenStudy (anonymous):

x-2+4=3 is x+2=3 which came to x=1 correct?

OpenStudy (anonymous):

|x-2|+4=3 subtract 4 from both sides

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so this would not have a solution

OpenStudy (anonymous):

because the answer I come up with is x=3 and

OpenStudy (anonymous):

Gj

OpenStudy (anonymous):

|x-6|=4 x-6=4 or x-6=-4 x=4+6 or x=-4+6 x=10 or x=2